Could you please confirm that I have configured the following formula for the length of an organ pipe correctly:
Given formula F = v/4(l + c ) where F is frequency, v is velocity of sound, l is pipe length and c is an end correction. Is it correct to say:
l + c = v/4f, that is is the correction is added to the length of the pipe or subtracted?

Answer by stanbon(75887)

F = (v/4)(l + c )
Multiply both sides by 4/v to get:
----
l+c = (4/v)F
l = (4F/v) - c
l = (4F-cv)/v
=================
Cheers,
Stan H.